# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

### v1

@@ Abstract
Given two populations from which independent binary observations are taken with parameters $p_{1}$ and $p_{2}$ respectively, estimators are proposed for the relative risk $p_{1}/p_{2}$, the odds ratio $p_{1}(1-p_{2})/(p_{2}(1-p_{1}))$ and their logarithms. The estimators guarantee that the relative mean-square error, or the mean-square error for the logarithmic versions, is less than a target value for any $p_{1},p_{2}\in(0,1)$, and the ratio of average sample sizes from the two populations is close to a prescribed value. The estimators can also be used with group sampling, whereby samples are taken in batches of fixed size from the two populations. The efficiency of the estimators with respect to the Cramér–Rao bound is good, and in particular it is close to $1$ for small values of the target error.
@@ 1 Introduction
When estimating any of these parameters, it is crucial to have knowledge about the accuracy of the estimation. Ideally, the estimation should guarantee a target accuracy, regardless of the unknown $p_{1}$ and $p_{2}$. A common measure of accuracy is the mean-square error (MSE), or its square root, known as root-mean-square error (RMSE), which for unbiased estimators reduce to variance or standard deviation respectively. For non-logarithmic parameters such as RR and OR, these error measurements are meaningful in a relative sense [16], because the significance of a given estimation error can only be assessed by comparing it with the true value of the parameter. Thus, it is natural to require that the RMSE be proportional to the true value of the parameter. On the other hand, for LRR and LOR the estimation error is meaningful by itself, without comparing with the true value, as the logarithm transforms multiplication into addition.
A second desirable feature, along with guaranteed accuracy, is to have control on the proportion of sample sizes of the two populations. Consider first the case that the two populations are sampled separately, meaning that samples from one or the other population can be taken as needed. In many use cases, it may be required that the two sample sizes be similar, or that they approximately satisfy a given ratio. Another possible sampling procedure is group sampling, whereby samples are collected in groups or batches of $l_{1}$ and $l_{2}$ samples from the two populations respectively. This imposes a strict sample size ratio of $l_{1}/l_{2}$. Either with separate sampling or with group sampling, these conditions only refer to the sample size ratio; the actual sample sizes should be chosen to fulfill the target accuracy.
@@
This paper presents unbiased estimators of the four parameters RR, LRR, OR and LOR, that guarantee a target accuracy and provide control on the sample size ratio, irrespective of $p_{1}$ and $p_{2}$. The estimators are based on the two-stage sampling procedure suggested in a previous work [16, section 4]. As argued previously, the target accuracy, $A$, is defined as relative MSE for RR and OR, or as MSE for LRR and LOR. The control on the sample size ratio means that, for separate sampling of each population, average sample sizes will approximately satisfy a specified proportion. The estimators can also be applied with group sequential sampling, using groups of $l_{1}$ and $l_{2}$ samples from each population. This incurs, as will be seen, a small increase in the average sample size compared to separate sampling, but ensures an exact ratio of the (random) sample sizes. In both cases the estimation efficiency, defined in terms of the Cramér–Rao bound, is considerably large, and is close to $1$ for small $A$.
@@
$I(x;u,v)=\frac{1}{\mathrm{B}(u,v)}\int_{0}^{x}t^{u-1}(1-t)^{v-1}\,\mathrm{d}t,$ (4)
@@
$f(y;u,v)=\frac{y^{u-1}}{\mathrm{B}(u,v)(1+y)^{u+v}},\quad y>0.$ (9)
@@ 2 Estimation of relative risk
The general estimation procedure for RR is motivated and described in Subsection 2.1, considering separate sampling of the two populations. Subsections 2.22.4 complete the definition of the estimator, obtain theoretical bounds for the average sample sizes and estimation efficiency, and compare them with simulation results. Subsection 2.5 describes the estimation method with group sampling, and then Subsections 2.6 and 2.7 obtain approximations for the average number of groups and efficiency in this case, and compare with simulation results.
2.1 Separate sampling. Estimation procedure
The estimator to be presented is unbiased, and uses two sampling stages. Each stage is comprised of two independent IBS procedures, one for each population. Before explaining the purpose of each stage, it is necessary to define several parameters and variables. The IBS parameters of the first stage are denoted as $r_{1}$ and $r_{2}$ for the two populations respectively. The resulting numbers of samples are negative binomial random variables, $M_{1}$ and $M_{2}$. Similarly, the second-stage IBS procedures have parameters $s_{1}$ and $s_{2}$, and the numbers of samples are $N_{1}$ and $N_{2}$. The total number of samples used from each population $i=1,2$ is then $M_{i}+N_{i}$, and from (11) it stems that
$\operatorname{E}[M_{i}+N_{i}]=\frac{r_{i}+s_{i}}{p_{i}}.$ (15)
@@
$\frac{\operatorname{E}[M_{1}+N_{1}]}{\operatorname{E}[M_{2}+N_{2}]}=\frac{r_{1}+s_{1}}{(r_{2}+s_{2})\theta}.$ (18)
@@
An initial idea to obtain information about $\theta$ from the first-stage variables $M_{1}$ and $M_{2}$ is to compute an estimate of this parameter using a generalization of the method described in Mendo [16, section 2] for estimating $p/(1-p)$. Namely, from (11) and (12) it follows that $(r_{1}-1)/(M_{1}-1)$ is an unbiased estimator of $p_{1}$ and $M_{2}/r_{2}$ is an unbiased estimator of $1/p_{2}$. Therefore, since the observations used by those estimators are independent,
@@
is an unbiased estimator of $\theta$. Replacing $\theta$ in (18) by this estimate suggests that the condition (17) will be roughly satisfied if $s_{1}$ and $s_{2}$ are chosen so that
@@
This will be referred to as error function. The parameters $\mu_{1},\mu_{2},\mu_{12}$ are introduced for convenience; this way the expression of $e(s_{1},s_{2})$ for other estimators will be the same as (29), only with different values of these parameters. Then, requiring
@@
2.2 Estimator parameters. Approximate average sample sizes
According to (15), to reduce the average sample sizes, $r_{i}$ and $s_{i}$, $i=1,2$ should be as small as possible. In this subsection approximate expressions are first derived for the average sample sizes, from which the choice of $s_{1}$, $s_{2}$ is addressed and the values of the estimator parameters $\gamma$, $\delta_{1}$, $\delta_{2}$, $\epsilon_{1}$, $\epsilon_{2}$ are determined. (Subsection 2.3 will discuss how to select the values of $r_{1}$, $r_{2}$.)
@@
Figure 1: Solutions $s_{1},s_{2}$ to (20) and (33) (example with $A=0.05$, $\delta_{1}=1$, $\delta_{2}=3$, $X=2.5$, $\mu_{1}=2$, $\mu_{2}=0$, $\mu_{12}=1$)
As indicated in Subsection 2.1, the solutions (34) and (35) have to be rounded, because only integer numbers can be used as IBS parameters. Depending on their specific values it may be necessary to round both of them up, or it may be sufficient to round one up and the other down, if that satisfies (31). A simple criterion, which will be assumed in the rest of the paper, is as follows. First, randomly choose $(i,j)=(1,2)$ or $(2,1)$ with equal probability; then round $s_{i}$ up and $s_{j}$ down, and check if those values satisfy (31). If not, try rounding $s_{i}$ down and $s_{j}$ up. If not valid either, round both values up, which necessarily satisfies (31).
@@
Figure 2 represents these approximations using the example values (same as in Figure 1) $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, and with $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ given by (30), as specified for RR. The figure illustrates that the accuracy of the approximations depends on the curvature of $s_{1}$ and $s_{2}$ as functions of $X$ and $1/X$. In fact, for certain values of $A$, $\delta_{1}$ and $\delta_{2}$ the approximations (37) and (38) are exact. This will be analyzed in Subsection 2.3.
@@
Figure 2: First-order approximations of $s_{1}$ and $s_{2}$ for RR (example with $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$)
@@
The term $\xi$ models the average increase in $s_{1}$ and $s_{2}$ due to rounding. The impact of this on $\operatorname{E}[M_{i}+N_{1}]$, $i=1,2$ will usually be negligible, and thus $\xi$ could be taken as 00 with good approximation. However, choosing $\xi=1$ (together with appropriate values of $r_{1}$ and $r_{2}$) will be useful in Subsection 2.3 to obtain upper bounds on $\operatorname{E}[M_{i}+N_{i}]$.
@@
The parameters $\mu_{1}$, $\mu_{2}$ $\mu_{12}$ for the RR estimator, given by (30), satisfy
@@
It is convenient to use a normalized version of the average sample sizes, $\operatorname{E}[M_{i}+N_{i}]\sqrt{p_{1}p_{2}}$, so that the normalized average sample sizes depend on the two probabilities only through their ratio $\theta$. Defining
@@
The average number of samples $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ is the sum of two terms inversely proportional to $p_{1}$ and $p_{2}$, according to (67) and (68); and it is $\lambda$ times more sensitive to $p_{2}$ than to $p_{1}$. That is, the parameter of the population for which a smaller average sample size is desired (as specified by (17)) has a stronger influence on both average sample sizes. It is also noteworthy that, for $\lambda$ fixed, (70) and (71) are minimized when $\lambda\theta=1$, which according to (17) and (18) corresponds to $r_{1}+s_{1}\approx r_{2}+s_{2}$. Thus, disparity in the sum of IBS parameters between the two populations incurs an increase in average sample sizes.
2.3 Analysis of the approximation. Bounds on average sample sizes
Based on an analysis of the approximation error in (37) and (38), this subsection obtains bounds on the average sample sizes and selects appropriate values for $r_{1}$ and $r_{2}$.
@@
The equality (75) with negative sign has the solution $A=-1/\mu_{12}$ if $\mu_{12}\neq 0$. With the estimators considered in this paper, $\mu_{12}$ is either 00 or positive, and this results in no solution or a negative solution for $A$, which is not valid. On the other hand, (75) with positive sign is satisfied if and only if $A(\delta_{1}+\mu_{1})^{2}+2(\delta_{1}+\mu_{1})-\mu_{12}=0.$ Taking into account (62), this can be expressed as $c(A,r_{1},\xi)=0$ with
@@
From (50) and (51) it is clear that $E[X]<r_{2}\theta/(r_{1}-1)$ and $\operatorname{E}[1/X]<r_{1}/((r_{2}-1)\theta)$. Substituting into (77) and (78) and using (39)–(42) and (60)–(62) yields
@@
The condition $c(A,r_{1},\xi)=0$, with $c(A,r_{1},\xi)$ given by (76), can be written as

### v2

@@ Abstract
Given two populations from which independent binary observations are taken with parameters $p_{1}$ and $p_{2}$ respectively, estimators are proposed for the relative risk $p_{1}/p_{2}$, the odds ratio $p_{1}(1-p_{2})/(p_{2}(1-p_{1}))$ and their logarithms. The sampling strategy used by the estimators is based on two-stage sequential sampling applied to each population, where the sample sizes of the second stage are computed from the results observed in the first stage. The estimators guarantee that the relative mean-square error, or the mean-square error for the logarithmic versions, is less than a target value for any $p_{1},p_{2}\in(0,1)$, and the ratio of average sample sizes from the two populations is close to a prescribed value. The estimators can also be used with group sampling, whereby samples are taken in batches of fixed size from the two populations simultaneously, each batch containing samples from the two populations. The efficiency of the estimators with respect to the Cramér–Rao bound is good, and in particular it is close to $1$ for small values of the target error.
@@ 1 Introduction
When estimating any of these parameters, it is crucial to have knowledge about the accuracy of the estimation. Ideally, the estimation should guarantee a target accuracy, regardless of the unknown $p_{1}$ and $p_{2}$. A common measure of accuracy is the mean-square error (MSE), or its square root, known as root-mean-square error (RMSE), which for unbiased estimators reduce to variance or standard deviation respectively. For non-logarithmic parameters such as RR and OR, these error measurements are meaningful in a relative sense [16], because the significance of a given estimation error can only be assessed by comparing it with the true value of the parameter. Thus, it is natural to require that the RMSE be proportional to the true value of the parameter. On the other hand, for LRR and LOR the estimation error is meaningful by itself, without comparing with the true value, as the logarithm transforms ratios into differences.
A second desirable feature, along with guaranteed accuracy, is to have control on the proportion of sample sizes of the two populations. Consider first the case that the two populations are sampled individually, i.e. with element sampling, meaning that samples from any population can be taken one by one as needed. In many use cases, it may be required that the two sample sizes be similar, or that they approximately satisfy a given ratio. Another possible sampling procedure is group sampling, whereby samples are collected in groups or batches, each containing $l_{1}$ samples from population 11 and $l_{2}$ from population $2$. This imposes a strict sample size ratio of $l_{1}/l_{2}$. Either with element sampling or with group sampling, these conditions only refer to the sample size ratio; the actual sample sizes should be chosen to fulfill the target accuracy.
Note that the difference of the group sampling scheme with respect to element sampling is not only that samples are taken in groups (as opposed to one by one), but also that each group simultaneously contains samples from both populations (instead of each population being sampled separately).
@@
This paper presents unbiased estimators of the four parameters RR, LRR, OR and LOR, that guarantee a target accuracy and provide control on the sample size ratio, irrespective of $p_{1}$ and $p_{2}$. The estimators are based on the two-stage sampling procedure suggested in a previous work [16, section 4]. As argued above, the target accuracy, $A$, is defined as relative MSE for RR and OR, or as MSE for LRR and LOR. The control on the sample size ratio means that, for element sampling of each population, average sample sizes will approximately satisfy a specified proportion. The estimators can also be applied with group sequential sampling, using groups of $l_{1}$ and $l_{2}$ samples from each population. This incurs, as will be seen, a small increase in the average sample size compared to element sampling, but ensures an exact ratio of the (random) sample sizes. In both cases the estimation efficiency, defined in terms of the Cramér–Rao bound, is considerably large, and is close to $1$ for small $A$.
@@
$I(x;u,v)=\frac{1}{\mathrm{B}(u,v)}\int_{0}^{x}t^{u-1}(1-t)^{v-1}\,\mathrm{d}t,\quad 0<x<1;\ u,v>0,$ (4)
@@
$f(y;u,v)=\frac{y^{u-1}}{\mathrm{B}(u,v)(1+y)^{u+v}},\quad y>0;\ u,v>0.$ (9)
@@ 2 Estimation of relative risk
The estimation procedure for RR with element sampling of the two populations is considered in Section 2.1. First, the general estimation approach is motivated and described (Section 2.1.1). The precise definition of the estimator is then completed (Section 2.1.2). Lastly, theoretical bounds are obtained for the average sample sizes and estimation efficiency, and these are compared with simulation results (Sections 2.1.3 and 2.1.4).
Group sampling is addressed in Section 2.2. The estimation procedure used in this case is presented (Section 2.2.1), and then approximations for the average number of groups and efficiency are obtained and compared with simulation results (Sections 2.2.2 and 2.2.3).
2.1 Element sampling
2.1.1 Estimation procedure
The estimator to be presented is unbiased, and uses two-stage sampling. Each of the two sampling stages is comprised of two independent IBS procedures, one for each population. Before explaining the purpose of each stage, it is necessary to define several parameters and variables. The IBS parameters of the first stage are denoted as $r_{1}$ and $r_{2}$ for the two populations respectively. The resulting numbers of samples are negative binomial random variables, $M_{1}$ and $M_{2}$. Similarly, the second-stage IBS procedures have parameters $s_{1}$ and $s_{2}$, and the numbers of samples are $N_{1}$ and $N_{2}$. The parameters $s_{1}$ and $s_{2}$ are obtained from $M_{1}$ and $M_{2}$, as will be seen, and are thus random variables. The total number of samples used from each population $i=1,2$ is then $M_{i}+N_{i}$. From (11) it stems that $\operatorname{E}[M_{i}]=r_{i}/p_{1}$, $\operatorname{E}[N_{i}\mid s_{i}]=s_{i}/p_{i}$, and
$\operatorname{E}[M_{i}+N_{i}]=\frac{r_{i}+\operatorname{E}[s_{i}]}{p_{i}}.$ (15)
@@ 2.1.1 Estimation procedure
$\frac{\operatorname{E}[M_{1}+N_{1}]}{\operatorname{E}[M_{2}+N_{2}]}=\frac{r_{1}+\operatorname{E}[s_{1}]}{(r_{2}+\operatorname{E}[s_{2}])\theta}.$ (18)
@@
An initial idea to obtain information about $\theta$ from the first-stage variables $M_{1}$ and $M_{2}$ is to compute an estimate of this parameter using a generalization of the method described in Mendo [16, section 2] for estimating p/(1p)p/(1-p). Namely, from (11) and (12) it follows that $(r_{1}-1)/(M_{1}-1)$ is an unbiased estimator of $p_{1}$ and $M_{2}/r_{2}$ is an unbiased estimator of $1/p_{2}$. Therefore, since the observations used by those estimators are independent,
@@
is an unbiased estimator of $\theta$. Replacing $\theta$ by this estimate and $\operatorname{E}[s_{i}]$ by $s_{i}$, $i=1,2$ in (18) suggests that the condition (17) will be roughly satisfied if $s_{1}$ and $s_{2}$ are chosen so that
@@
This will be referred to as error function. The parameters $\mu_{1}$, $\mu_{2}$ and $\mu_{12}$ are introduced for convenience; this way the expression of $e(s_{1},s_{2})$ for other estimators will be the same as (29), only with different values of these parameters. Then, requiring
@@
2.1.2 Estimator parameters and approximate average sample sizes
According to (15), to achieve small average sample sizes, $r_{i}$ and $s_{i}$, $i=1,2$ should be as small as possible. In this section approximate expressions are first derived for the average sample sizes, from which the choice of $s_{1}$, $s_{2}$ is addressed and the values of the estimator parameters $\gamma$, $\delta_{1}$, $\delta_{2}$, $\epsilon_{1}$, $\epsilon_{2}$ are determined. (Section 2.1.3 will discuss how to select the values of $r_{1}$, $r_{2}$.)
@@ 2.1.2 Estimator parameters and approximate average sample sizes
Figure 1: Solutions $s_{1},s_{2}$ to (20) and (33) (example with $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, $X=2.5$, $\mu_{1}=2$, $\mu_{2}=0$, $\mu_{12}=1$)
As indicated in Section 2.1.1, the solutions (34) and (35) have to be rounded, because only integer numbers can be used as IBS parameters. Depending on their specific values it may be necessary to round both of them up, or it may be sufficient to round one up and the other down, if that satisfies (31). A simple criterion, which will be assumed in the rest of the paper, is as follows. First, randomly choose $(i,j)=(1,2)$ or $(2,1)$ with equal probability; then round $s_{i}$ up and $s_{j}$ down, and check if those values satisfy (31). If not, try rounding $s_{i}$ down and $s_{j}$ up. If not valid either, round both values up, which necessarily satisfies (31).
@@
Figure 2 represents these approximations using the example values (same as in Figure 1) $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, and with $\mu_{1}$, $\mu_{2}$, $\mu_{12}$ given by (30), as specified for RR. The figure illustrates that the accuracy of the approximations depends on the curvature of $s_{1}$ and $s_{2}$ as functions of $X$ and $1/X$. In fact, for certain values of $A$, $\delta_{1}$ and $\delta_{2}$ the approximations (37) and (38) are exact. This will be analyzed in Section 2.1.3.
@@
Figure 2: First-order approximations of $s_{1}$ and $s_{2}$ for RR (example with $A=0.05$, $\gamma=0.5$, $\delta_{1}=1$, $\delta_{2}=3$, $\mu_{1}=2$, $\mu_{2}=0$, $\mu_{12}=1$)
@@
The term $\xi$ models the average increase in $s_{1}$ and $s_{2}$ due to rounding. The impact of this on $\operatorname{E}[M_{i}+N_{1}]$, $i=1,2$ will usually be negligible, and thus $\xi$ could be taken as 0 with good approximation. However, choosing $\xi=1$ (together with appropriate values of $r_{1}$ and $r_{2}$) will be useful in Section 2.1.3 to obtain upper bounds on $\operatorname{E}[M_{i}+N_{i}]$.
@@
The parameters $\mu_{1}$, $\mu_{2}$ and $\mu_{12}$ for the RR estimator, given by (30), satisfy
@@
It will be beneficial to use normalized versions of the average sample sizes, $\operatorname{E}[M_{i}+N_{i}]\sqrt{p_{1}p_{2}}$, so that they depend on the two probabilities only through their ratio $\theta$. Defining
@@
The average number of samples $\operatorname{E}[M_{i}+N_{i}]$, $i=1,2$ is the sum of two terms inversely proportional to $p_{1}$ and $p_{2}$, according to (67) and (68); and it is $\lambda$ times more sensitive to $p_{2}$ than to $p_{1}$. That is, the parameter of the population for which a smaller average sample size is desired (as specified by (17)) has a stronger influence on both average sample sizes. It is also noteworthy that, for $\lambda$ fixed, (70) and (71) are minimized when $\lambda\theta=1$, which according to (17) and (18) corresponds to $r_{1}+\operatorname{E}[s_{1}]\approx r_{2}+\operatorname{E}[s_{2}]$.
2.1.3 Analysis of the approximation and bounds on average sample sizes
Based on an analysis of the approximation error in (37) and (38), this section obtains bounds on the average sample sizes and selects appropriate values for $r_{1}$ and $r_{2}$.
@@ 2.1.3 Analysis of the approximation and bounds on average sample sizes
The equality (75) with negative sign has the solution $A=-1/\mu_{12}$ if $\mu_{12}\neq 0$. With the estimators considered in this paper, $\mu_{12}$ is either 0 or positive, and this results in no solution or a negative solution for $A$, which is not valid. On the other hand, (75) with positive sign is satisfied if and only if $A(\delta_{1}+\mu_{1})^{2}+2(\delta_{1}+\mu_{1})-\mu_{12}=0.$ Taking into account (62), this can be expressed as $c(A,r_{1},\xi)=0$ with
@@
From (50) and (51) it is clear that $\operatorname{E}[X]<r_{2}\theta/(r_{1}-1)$ and $\operatorname{E}[1/X]<r_{1}/((r_{2}-1)\theta)$. Substituting into (77) and (78) and using (39)–(42) and (60)–(62) yields
@@
The condition $c(A,r_{1},\xi)=0$ can be written, from (76), as